# Selmer stability for elliptic curves in Galois $\ell$-extensions

### v1

@@ Abstract.
We study the behavior of Selmer groups of an elliptic curve $E/\mathbb{Q}$ in finite Galois extensions with prescribed Galois group. Fix a prime $\ell\geq 5$, a finite group $G$ with #$\#G=\ell^{n}$, and an elliptic curve $E/\mathbb{Q}$ with $\operatorname{Sel}_{\ell}(E/\mathbb{Q})=0$ and surjective $-\ell$ Galois representation. We show that there exist infinitely many Galois extensions $F/\mathbb{Q}$ with Galois group $\operatorname{Gal}(F/\mathbb{Q})\simeq G$ for which the $\ell$-Selmer group $\operatorname{Sel}_{\ell}(E/F)$ also vanishes. We obtain an asymptotic lower bound for the number $\mathcal{M}(G,E;X)$ of such fields $F$ with absolute discriminant $|\Delta_{F}|\leq X$, proving that there is an explicit constant $\delta>0$ such that
@@ 1.1. Background and motivation
Let $E$ be an elliptic curve defined over a number field $K$. The group $E(K)$ of $K$-rational points, known as the Mordell–Weil group, is finitely generated, and its rank is a fundamental arithmetic invariant. One fruitful approach to studying the structure of $E(K)$ is through Galois cohomology. For a fixed integer $n\geq 1$, the $n$-torsion subgroup $E[n]\subset E(\overline{K})$ carries a natural action of the absolute Galois group $\operatorname{G}_{K}:=\operatorname{Gal}(\overline{K}/K)$. The Selmer group $\operatorname{Sel}_{n}(E/K)\subset H^{1}(K,E[n])$, defined by imposing local conditions at all primes of $K$, sits in a natural short exact sequence,
@@
In particular, these diophantine stability questions are studied for cyclic extensions of $\mathbb{Q}$ with prime degree. Mazur and Rubin [MR10] studied the distribution of 2-Selmer ranks in families of quadratic twists of a fixed elliptic curve. Their techniques yield results on the stability of rank across large sets of quadratic twists, and under certain conditions, also demonstrate that arbitrarily large ranks can arise after base change by quadratic extensions. For any prime $\ell$, Klagsbrun–Mazur–Rubin [KMR14], Mazur–Rubin [MRL18] have studied the behavior of $\ell$-Selmer groups in $\mathbb{Z}/\ell\mathbb{Z}$-extensions, establishing patterns of stability and growth. Smith has recently obtained similar results using arithmetic statistics and the geometry of numbers [Smi22, Smi23].
The broader phenomenon of rank growth and stability in families of number field extensions has attracted significant recent attention. For example, see the work of Lemke-Oliver and Thorne [LOT21], Shnidman and Weiss [SW23], Beneish–Kundu–Ray [BKR24], Berg–Ryan–Young [BRY24], Keliher [Kel24], Pathak–Ray [PR25] and Keliher–Park [KP25].
Understanding Galois extensions of a fixed number field with a prescribed Galois group is a fundamental problem in number theory. Conjectures predicting the asymptotic number of such extensions, ordered by their absolute discriminant, were proposed by Malle [Mal02a, Mal04]. These conjectures may be viewed as statistical refinements of the classical inverse Galois problem. Substantial progress towards this has been made in special cases. For example, the conjecture is known for abelian groups, due to the work of Mäki [Mäk85] and Wright [Wri89], and more recently for certain nilpotent groups by Koymans and Pagano [KP23]. Although the inverse Galois problem for finite solvable groups over number fields was resolved by Shafarevich [Š54], Malle’s conjecture remains open for general solvable groups.
@@ 1.2. Main results
$\rho_{E,\ell}:\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\longrightarrow\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$
@@
Let $\mathcal{G}$ be a finite group and $\pi_{\mathcal{G}}:\mathcal{G}\hookrightarrow S_{\#\mathcal{G}}$ denote the regular representation of $\mathcal{G}$. For $g\in\mathcal{G}$, define the index of $g$, denoted $\operatorname{ind}(g)$, as
$\operatorname{ind}(g):=\#\mathcal{G}-\#\text{ orbits of }\pi_{\mathcal{G}}(g).$
We set $a(\mathcal{G}):=\left(\min_{g\neq 1}\operatorname{ind}(g)\right)^{-1}$ and note that when $\mathcal{G}$ is an $\ell$-group with $\ell^{n}$ elements, $a(\mathcal{G})=\left(\ell^{n-1}(\ell-1)\right)^{-1}$. The weak form of Malle’s conjecture (cf. [Mal02b, p.316]) predicts that for any $\epsilon>0$, there exist constants $c_{1}(\mathcal{G}),c_{2}(\mathcal{G};\epsilon)>0$ such that
@@
for all sufficiently large $X$. This has been established for all finite nilpotent groups by Klüners and Malle [KM04]. For finite nilpotent groups satisfying additional hypotheses, Koymans and Pagano [KP23] have proven a stronger version of Malle’s conjecture, obtaining the exact asymptotic.
On the other hand, let $\mathcal{M}(\mathcal{G},E;X)$ denote the number of Galois extensions $F/\mathbb{Q}$ with $\operatorname{Gal}(F/\mathbb{Q})\simeq\mathcal{G}$, $|\Delta_{F}|\leq X$, and $\operatorname{Sel}_{\ell}(E/F)=0$. Note that for such fields $F/\mathbb{Q}$, $\operatorname{rank}E(F)=0$ and $\Sh(E/F)[\ell]=0$. For a group $G$ with #$\#G=\ell^{n}$, our main result establishes asymptotic lower bounds for $\mathcal{M}(G,E;X)$ as $x\to\infty$, showing that it grows comparably to $\mathcal{N}(G;X)$. This demonstrates that the vanishing of the $\ell$-Selmer group over such $G$-extensions is not a rare phenomenon.
@@
where $\delta:=\frac{\ell^{2}-\ell-1}{\ell^{n-1}(\ell^{2}-1)}$.
@@
For $\ell=5$, it is proven by Bhargava and Shankar [BS13] that there is a positive density of elliptic curves (ordered by height) $E_{/\mathbb{Q}}$ for which $\operatorname{Sel}_{\ell}(E/\mathbb{Q})=0$. On the other hand, a result of Duke [Duk97] shows that $\rho_{E,\ell}$ is surjective for almost all elliptic curves $E_{/\mathbb{Q}}$. Thus for $\ell=5$, the conditions of Theorem 1.1 hold for a positive density of elliptic curves over $\mathbb{Q}$.
@@ 1.3. Methodology
holds. The argument then follows via induction on the length of $G$. The inverse Galois problem for finite $\ell$-groups is known from the work of Reichardt [Rei37] and Scholz [Sch37]. The Galois cohomological strategy allows us to inductively construct towers of number fields
@@ 1.4. Organization
Including the introduction, the article is organized into four sections. Section 2 revisits the classical strategy of Reichardt [Rei37] and Scholz [Sch37] for resolving the inverse Galois problem for finite $\ell$-groups over $\mathbb{Q}$. Our emphasis is on making explicit which primes split or ramify in such extensions. This information is essential for analyzing the behavior of Selmer groups in Galois towers. In Section 3, we establish sufficient conditions on an $\ell$-cyclic extension $L/K$ under which the vanishing of $\operatorname{Sel}_{\ell}(E/K)$ implies the vanishing of $\operatorname{Sel}_{\ell}(E/L)$. These build upon previous constructions in [PR25, Section 2]. We conclude this section by proving the existence of infinitely many $G$-extensions $L/\mathbb{Q}$ such that $\operatorname{Sel}_{\ell}(E/L)=0$. Finally, Section 4 is devoted to the proof of Theorem 1.1. We recall key results of Klüners and Malle concerning the enumeration of $G$-extensions of $\mathbb{Q}$ ordered by absolute discriminant. The problem is reduced to counting global Galois cohomology classes subject to prescribed local conditions. To estimate the size of these cohomology groups, we invoke a formula of Wiles which yields precise asymptotics, allowing us to deduce the desired lower bounds.
@@ 2. The inverse Galois problem for finite $\ell$-groups
Shafarevich proved that given any number field $K$ and a finite solvable group $G$, there exists a Galois extension $L/K$ with $\operatorname{Gal}(L/K)\simeq G$ (see [Saf54] and [NSW08, Chapter IX, Section 6]). Let $\ell$ be an odd prime number. Given a natural number $k$, let $\mu_{k}$ denote the set of $k$-th roots of unity. Fix throughout this article a finite group $G$ with #$\#G=\ell^{n}$. In this special case, the inverse Galois problem over $\mathbb{Q}$ for the group $G$ is due to Reichardt [Rei37] and Scholz [Sch37]. An exposition of their method can be found in [Mas06]. In this section, we review this construction with a view towards its application in establishing diophantine stability results, for which we need an understanding of the ramification of primes. More precisely, we shall see that these sets of primes are prescribed by Chebotarev conditions.
@@ 2.1. The embedding problem
[…]
Let $\operatorname{G}_{\mathbb{Q}}:=\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ and assume that there exists a Galois extension $L/\mathbb{Q}$ such that ${\operatorname{Gal}(L/\mathbb{Q})\simeq G}$. This gives rise to a surjection $\varphi:\operatorname{G}_{\mathbb{Q}}\rightarrow G$ such that ¯$\overline{\mathbb{Q}}^{\operatorname{ker}\varphi}=L$. The embedding problem asks whether $L$ can be embedded into a larger Galois extension $\widetilde{L}/\mathbb{Q}$ along with an isomorphism $\operatorname{Gal}(\widetilde{L}/\mathbb{Q})\simeq\widetilde{G}$, such that the following diagram commutes:
@@
In the above diagram, the downward arrows are isomorphisms. Equivalently, the embedding problem is solvable if $\varphi$ can be lifted to a surjective homomorphism $\widetilde{\varphi}:\operatorname{G}_{\mathbb{Q}}\twoheadrightarrow\widetilde{G}$. Indeed, setting $\widetilde{L}:=\overline{\mathbb{Q}}^{\operatorname{ker}\widetilde{\varphi}}$ we find that $\operatorname{Gal}(\widetilde{L}/\mathbb{Q})\simeq\widetilde{G}$. Note that if the extension (2.2) is non-split then any lift $\widetilde{\varphi}$ is surjective. In order to inductively construct the extensions $L_{i}$, it suffices to solve the embedding problem for
@@ 2.2. Cohomological parametrization of extension classes
[…]
Then $A$ has an induced $G$-module structure which we describe. Let $\eta:G\rightarrow\widetilde{G}$ be a set theoretic section of the map $\pi:\widetilde{G}\rightarrow G$. Since $\pi$ is surjective, such a map exists. We define the action of $G$ on $A$ as follows: for $a\in A$ and $\sigma\in G$,
$\sigma\cdot a:=\eta(\sigma)\,a\,\eta(\sigma)^{-1}.$
There are a few things to take note of here. First, since $A$ is a normal subgroup of $G$, the element $\eta(\sigma)\,a\,\eta(\sigma)^{-1}$ is in $A$. Next, since $A$ is abelian, $\sigma\cdot a$ is independent of the choice of set theoretic section $\eta$. From here on, when referring to $A$ as a $G$-module, it will be with respect to this chosen action, and $H^{i}(G,A)$ will be the associated cohomology groups. Two extensions $\widetilde{G}_{1}$ and $\widetilde{G}_{2}$ are said to be equivalent if there is an isomorphism $f:\widetilde{G}_{1}\xrightarrow{\sim}\widetilde{G}_{2}$ such that the following diagram commutes:
@@
[…]
    such that for any set theoretic section $\eta:G\rightarrow\widetilde{G}$ of $\pi$, we have that $\sigma\cdot a=\eta(\sigma)\,a\,\eta(\sigma)^{-1}$
@@
The class associated to $\widetilde{G}$ is denoted by $\theta_{\widetilde{G}}\in H^{2}(G,A)$. Moreover, the association has the property that the trivial class $\widetilde{G}:=G\ltimes A$ corresponds to the trivial element in $H^{2}(G,A)$.
@@
This is a standard result, cf. [Wei94, Theorem 6.6.3]. ∎

### v2

@@ Abstract.
We study the behavior of Selmer groups of an elliptic curve $E/\mathbb{Q}$ in finite Galois extensions with prescribed Galois group. Fix a prime $\ell\geq 5$, a finite group $G$ with $G=\ell^{n}$, and an elliptic curve $E/\mathbb{Q}$ with $\operatorname{Sel}_{\ell}(E/\mathbb{Q})=0$ and surjective mod-$\ell$ Galois representation. We show that there exist infinitely many Galois extensions $F/\mathbb{Q}$ with Galois group $\operatorname{Gal}(F/\mathbb{Q})\simeq G$ for which the $\ell$-Selmer group $\operatorname{Sel}_{\ell}(E/F)$ also vanishes. We obtain an asymptotic lower bound for the number $\mathcal{M}(G,E;X)$ of such fields $F$ with absolute discriminant $|\Delta_{F}|\leq X$, proving that there is an explicit constant $\delta>0$ such that
@@
2020 Mathematics Subject Classification:
11G05, 11R45
Research of the first author was partially supported by an INSPIRE faculty fellowship.
Data availability statement: Data sharing not applicable – no new data generated, or the article describes entirely theoretical research.
Conflict of interest statement: The authors do not have any conflict of interest to declare.
@@ 1.1. Background and motivation
Let $E$ be an elliptic curve defined over a number field $K$. The group $E(K)$ of $K$-rational points, known as the Mordell–Weil group, is finitely generated, and its rank is a fundamental arithmetic invariant. One fruitful approach to studying the structure of $E(K)$ is through Galois cohomology. For a fixed integer $n\geq 1$, the $n$-torsion subgroup $E[n]\subset E(\overline{K})$ carries a natural action of the absolute Galois group $\operatorname{G}_{K}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Gal}(\overline{K}/K)$. The Selmer group $\operatorname{Sel}_{n}(E/K)\subset H^{1}(K,E[n])$, defined by imposing local conditions at all primes of $K$, sits in a natural short exact sequence,
@@
In particular, these diophantine stability questions are studied for cyclic extensions of $\mathbb{Q}$ with prime degree. Mazur and Rubin [MazurRubinQtwists] studied the distribution of 2-Selmer ranks in families of quadratic twists of a fixed elliptic curve. Their techniques yield results on the stability of rank across large sets of quadratic twists, and under certain conditions, also demonstrate that arbitrarily large ranks can arise after base change by quadratic extensions. For any prime $\ell$, Klagsbrun–Mazur–Rubin [KMR], Mazur–Rubin [mazur2018diophantine] have studied the behavior of $\ell$-Selmer groups in $\mathbb{Z}/\ell\mathbb{Z}$-extensions, establishing patterns of stability and growth. Smith has recently obtained similar results using arithmetic statistics and the geometry of numbers [smith2022distribution1, smith2022distribution2].
The broader phenomenon of rank growth and stability in families of number field extensions has attracted significant recent attention. For example, see the work of Lemke-Oliver and Thorne [Oliverthorne], Shnidman and Weiss [ShnidmanWeiss], Beneish–Kundu–Ray [BKR], Berg–Ryan–Young [Berg], Keliher [keliher], Pathak–Ray [pathakray] and Keliher–Park [keliherpark].
Understanding Galois extensions of a fixed number field with a prescribed Galois group is a fundamental problem in number theory. Conjectures predicting the asymptotic number of such extensions, ordered by their absolute discriminant, were proposed by Malle [Malle1, Malle2]. These conjectures may be viewed as statistical refinements of the classical inverse Galois problem. Substantial progress towards this has been made in special cases. For example, the conjecture is known for abelian groups, due to the work of Mäki [maki1985density] and Wright [wright1989distribution], and more recently for certain nilpotent groups by Koymans and Pagano [koymans2023malle]. Although the inverse Galois problem for finite solvable groups over number fields was resolved by Shafarevich [Shaferevich], Malle’s conjecture remains open for general solvable groups.
@@ 1.2. Main results
$\rho_{E,\ell}\mathrel{\mathop{\ordinarycolon}}\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\longrightarrow\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$
@@
Let $\mathcal{G}$ be a finite group and $\pi_{\mathcal{G}}\mathrel{\mathop{\ordinarycolon}}\mathcal{G}\hookrightarrow S_{\#\mathcal{G}}$ denote the regular representation of $\mathcal{G}$. For $g\in\mathcal{G}$, define the index of $g$, denoted $\operatorname{ind}(g)$, as
$\operatorname{ind}(g)\mathrel{\mathop{\ordinarycolon}}=\#\mathcal{G}-\#\text{ orbits of }\pi_{\mathcal{G}}(g).$
We set $a(\mathcal{G})\mathrel{\mathop{\ordinarycolon}}=\left(\min_{g\neq 1}\operatorname{ind}(g)\right)^{-1}$ and note that when $\mathcal{G}$ is an $\ell$-group with $\ell^{n}$ elements, $a(\mathcal{G})=\left(\ell^{n-1}(\ell-1)\right)^{-1}$. The weak form of Malle’s conjecture (cf. [Malle1, p.316]) predicts that for any $\epsilon>0$, there exist constants $c_{1}(\mathcal{G}),c_{2}(\mathcal{G};\epsilon)>0$ such that
@@
for all sufficiently large $X$. This has been established for all finite nilpotent groups by Klüners and Malle [klunersmalle]. For finite nilpotent groups satisfying additional hypotheses, Koymans and Pagano [koymans2023malle] have proven a stronger version of Malle’s conjecture, obtaining the exact asymptotic.
On the other hand, let $\mathcal{M}(\mathcal{G},E;X)$ denote the number of Galois extensions $F/\mathbb{Q}$ with $\operatorname{Gal}(F/\mathbb{Q})\simeq\mathcal{G}$, $|\Delta_{F}|\leq X$, and $\operatorname{Sel}_{\ell}(E/F)=0$. Note that for such fields $F/\mathbb{Q}$, $\operatorname{rank}E(F)=0$ and $\Sh(E/F)[\ell]=0$. Thus, the Mordell–Weil rank and the $\ell$-part of the Tate-Sharevich group of $E$ remain stable in $F/\mathbb{Q}$. For a group $G$ with $G=\ell^{n}$, our main result establishes asymptotic lower bounds for $\mathcal{M}(G,E;X)$ as $x\to\infty$, showing that it grows comparably to $\mathcal{N}(G;X)$. This demonstrates that the vanishing of the $\ell$-Selmer group over such $G$-extensions is not a rare phenomenon.
@@
where $\delta\mathrel{\mathop{\ordinarycolon}}=\frac{\ell^{2}-\ell-1}{\ell^{n-1}(\ell^{2}-1)}$.
@@
For $\ell=5$, it is proven by Bhargava and Shankar [bhargavashankar5] that there is a positive density of elliptic curves (ordered by height) $E_{/\mathbb{Q}}$ for which $\operatorname{Sel}_{\ell}(E/\mathbb{Q})=0$. On the other hand, a result of Duke [duke] shows that $\rho_{E,\ell}$ is surjective for almost all elliptic curves $E_{/\mathbb{Q}}$. Thus for $\ell=5$, the conditions of Theorem 1.1 hold for a positive density of elliptic curves over $\mathbb{Q}$.
It is conceivable that the results of this article can be generalized to elliptic curves defined over arbitrary number fields. However, for ease of exposition, we restrict ourselves in this paper to the case of elliptic curves defined over $\mathbb{Q}$.
@@ 1.3. Methodology
holds. The argument then follows via induction on the length of $G$. The inverse Galois problem for finite $\ell$-groups is known from the work of Reichardt [Reichardt] and Scholz [Scholz]. The Galois cohomological strategy allows us to inductively construct towers of number fields
@@ 1.4. Organization
Including the introduction, the article is organized into four sections. Section 2 revisits the classical strategy of Reichardt [Reichardt] and Scholz [Scholz] for resolving the inverse Galois problem for finite $\ell$-groups over $\mathbb{Q}$. Our emphasis is on making explicit which primes split or ramify in such extensions. This information is essential for analyzing the behavior of Selmer groups in Galois towers. In Section 3, we establish sufficient conditions on an $\ell$-cyclic extension $L/K$ under which the vanishing of $\operatorname{Sel}_{\ell}(E/K)$ implies the vanishing of $\operatorname{Sel}_{\ell}(E/L)$. These build upon previous constructions in [pathakray, Section 2]. We conclude this section by proving the existence of infinitely many $G$-extensions $L/\mathbb{Q}$ such that $\operatorname{Sel}_{\ell}(E/L)=0$. Finally, Section 4 is devoted to the proof of Theorem 1.1. We recall key results of Klüners and Malle concerning the enumeration of $G$-extensions of $\mathbb{Q}$ ordered by absolute discriminant. The problem is reduced to counting global Galois cohomology classes subject to prescribed local conditions. To estimate the size of these cohomology groups, we invoke a formula of Wiles which yields precise asymptotics, allowing us to deduce the desired lower bounds.
@@ 2. The inverse Galois problem for finite $\ell$-groups
Shafarevich proved that given any number field $K$ and a finite solvable group $G$, there exists a Galois extension $L/K$ with $\operatorname{Gal}(L/K)\simeq G$ (see [Shafarevich] and [NSW, Chapter IX, Section 6]). Let $\ell$ be an odd prime number. Given a natural number $k$, let $\mu_{k}$ denote the set of $k$-th roots of unity. Fix throughout this article a finite group $G$ with $G=\ell^{n}$. In this special case, the inverse Galois problem over $\mathbb{Q}$ for the group $G$ is due to Reichardt [Reichardt] and Scholz [Scholz]. An exposition of their method can be found in [massey2006inverse]. In this section, we review this construction with a view towards its application in establishing diophantine stability results, for which we need an understanding of the ramification of primes. More precisely, we shall see that these sets of primes are prescribed by Chebotarev conditions.
@@ 2.1. The embedding problem
$1\to\mathbb{Z}/\ell\mathbb{Z}\xrightarrow{\iota}\widetilde{G}\xrightarrow{\pi}G\rightarrow 1.$ (2.2)
Let $\operatorname{G}_{\mathbb{Q}}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ and assume that there exists a Galois extension $L/\mathbb{Q}$ such that $\operatorname{Gal}(L/\mathbb{Q})\simeq G$}. This gives rise to a surjection $\varphi\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow G$ such that $\overline{\mathbb{Q}}^{\operatorname{ker}\varphi}=L$. The embedding problem asks whether $L$ can be embedded into a larger Galois extension $\widetilde{L}$$\mathbb{Q}$ along with an isomorphism $\operatorname{Gal}(\widetilde{L}/\mathbb{Q})\simeq\widetilde{G}$, such that the following diagram commutes:
@@
In the above diagram, the downward arrows are isomorphisms. Equivalently, the embedding problem is solvable if $\varphi$ can be lifted to a surjective homomorphism $\widetilde{\varphi}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\twoheadrightarrow\widetilde{G}$. Indeed, setting $\widetilde{L}\mathrel{\mathop{\ordinarycolon}}=\overline{\mathbb{Q}}^{\operatorname{ker}\widetilde{\varphi}}$ we find that $\operatorname{Gal}(\widetilde{L}/\mathbb{Q})\simeq\widetilde{G}$. Note that if the extension (2.2) is non-split then any lift $\widetilde{\varphi}$ is surjective. In order to inductively construct the extensions $L_{i}$, it suffices to solve the embedding problem for
@@ 2.2. Cohomological parametrization of extension classes
$1\rightarrow A\rightarrow\widetilde{G}\xrightarrow{\pi}G\rightarrow 1.$
Then $A$ has an induced $G$-module structure which we describe. Let $\eta\mathrel{\mathop{\ordinarycolon}}G\rightarrow\widetilde{G}$ be a set theoretic section of the map $\pi\mathrel{\mathop{\ordinarycolon}}\widetilde{G}\rightarrow G$. Since $\pi$ is surjective, such a map exists. We define the action of $G$ on $A$ as follows: for $a\in A$ and $\sigma\in G$,
$\sigma\cdot a\mathrel{\mathop{\ordinarycolon}}=\eta(\sigma)\,a\,\eta(\sigma)^{-1}.$
There are a few things to take note of here. First, since $A$ is a normal subgroup of $G$, the element $\eta(\sigma)\,a\,\eta(\sigma)^{-1}$ is in $A$. Next, since $A$ is abelian, $\sigma\cdot a$ is independent of the choice of set theoretic section $\eta$. From here on, when referring to $A$ as a $G$-module, it will be with respect to this chosen action, and $H^{i}(G,A)$ will be the associated cohomology groups. Two extensions $\widetilde{G}_{1}$ and $\widetilde{G}_{2}$ are said to be equivalent if there is an isomorphism $f\mathrel{\mathop{\ordinarycolon}}\widetilde{G}_{1}\xrightarrow{\sim}\widetilde{G}_{2}$ such that the following diagram commutes:
@@
    $1\mapsto A\xrightarrow{\iota}\widetilde{G}\xrightarrow{\pi}G\rightarrow 1$
    such that for any set theoretic section $\eta\mathrel{\mathop{\ordinarycolon}}G\rightarrow\widetilde{G}$ of $\pi$, we have that $\sigma\cdot a=\eta(\sigma)\,a\,\eta(\sigma)^{-1}$
@@
The class associated to $\widetilde{G}$ is denoted by $\theta_{\widetilde{G}}\in H^{2}(G,A)$. Moreover, the association has the property that the trivial class $\widetilde{G}\mathrel{\mathop{\ordinarycolon}}=G\ltimes A$ corresponds to the trivial element in $H^{2}(G,A)$.
@@
This is a standard result, cf. [weibel, Theorem 6.6.3]. ∎